hiiiihiiiiiiiiii:

Subtract this problem using mixed numbers. Reduce your answer to the lowest term. 16 3/4 - 8 5/12 8 4/12 8 1/3 8 8 2/8 I thanks its D

hiiiihiiiiiiiiii:

isent it D

hiiiihiiiiiiiiii:

right

ramen:

I'll take you through the steps: 1. 16-8=8 for the number that isn't a fraction. 2. Find the common denominator of 3/4 and 5/12. So 4 can go into 12 three times so 4*3=12 and 3*3=9. We now have the fractions 9/12 and 5/12. 3. Subtract 5/12 from 9/12. 9-5=4 and the denominator stays the same so we have 4/12. What is the answer?

hiiiihiiiiiiiiii:

D

ramen:

Does 4/12 = 2/8?

hiiiihiiiiiiiiii:

nooooooooo

ramen:

So what is the correct answer then?

hiiiihiiiiiiiiii:

A

ramen:

Yes that is correct. Although 4/12 can be simplified to 1/3, that isn't in our answer choices.

hiiiihiiiiiiiiii:

ok ty

ramen:

np.